31. Find the common difference of the arithmetic sequence {by} = {3n + 1} and the first four terms.
Answer #1

Giren arithmatic sequence NASA < bm} = { snti} Common difference = 3 = Coethriest of - bo= ba = 3(1+14 b2=3(2) +17 63 = 3(3)4